Part III: Set up the circuit on a breadboard, use R1 = 101, R2 512, and R3 41, and adjust the DC power supply to Vs,=10V and1 12 R R2 +VRI- +VR2- Iz Vsi R3 M +VR3- Vs2

Can it be shown on paper please
Answer #1

Given that 1 12 Rocion Ov R2:50 um W + - VR 1,12 Ver ti Vs R₂ SV lov بها Apply nodal modal analysis at mode -o I = Izt I3 Vs,Another method - 0 Apply Kill to loop - 0 +10 + II R + Iz R3 = 0 I3 = I, I2 10 + II RI + (1, -12) R3 = 0 - 10 IR + I, R3 – I₂IUT, - 412 - 41 +91) O x2 28 JI - 20 - 281 -812 +6312 551) -35 = -55 CS Scanned with CamScannerT2 Substitute in Iz equation -0 141, - 412 - - 10 1411 tu - 10 1411 -14 I -IA Iz = I2 - Il = OĄ VRI = ,I,xri -1X10 -10 V VR2